1. Find the arc length of the graph of the function over the indicated interval. (x2 + 2)3/2 3 -;[0,1]
Answer #1

YE (0+2.9312 [0,1] length = 1+ en die Home 2 (2+2) (201) dx = 4(+2)!! Cength= 11+2 (2²+2) du - *+232 +day = [(x+1)? du - l